OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (−4, −five fourths) and a directrix of y = twenty seven fourths? f(x) =negative one sixteenthsx2 − one half x + seven-fourths f(x) =negative one sixteenthsx2 − one half x + thirteen fourths f(x) =one sixteenthsx2 + one half x + seven-fourths f(x) =one sixteenthsx2 + one half x + thirteen fourths CAN U PLEASE HELP ME? @hartnn

OpenStudy (anonymous):

hello? @hartnn

hartnn (hartnn):

solved any similar problem before ?

OpenStudy (anonymous):

nope

hartnn (hartnn):

see the example we can use similar approach for this problem

hartnn (hartnn):

Distance between (x,y) and (-4,-4/5) is ??

OpenStudy (anonymous):

this is so hard

hartnn (hartnn):

if you know the distance formula, then it isn't

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

hartnn (hartnn):

so once again Distance between (x,y) and (-4,-4/5) is ??

OpenStudy (anonymous):

where is x^2?

hartnn (hartnn):

which x^2 ? use the points and the formula i gave to find the distance x1 = x y1 =y x2 =-4 y2= -4/5

OpenStudy (anonymous):

ooohh ok

OpenStudy (anonymous):

\[\sqrt{-4}-(-4)^{2} +(\frac{ -5 }{ 4}-(-\frac{ 4 }{ 5})^{2}\]

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

where is x and y ? :O

hartnn (hartnn):